# FLLIC: Functionally Lossless Image Compression

### v1

@@ Abstract
I Introduction
Accompanying the exciting progress of modern machine learning with deep neural networks (DNNs), many researchers have published a family of end-to-end optimized DNN image compression methods in the past five years. Most of these methods are rate-distortion optimized for lossy compression  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. By design, they cannot perform lossless or near-lossless image compression even with an unlimited bit budget. More recently, a number of research teams embark on developing DNN lossless image compression methods, aiming at minimum code length [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. These authors apply various deep neural networks, such as autoregressive models [33, 34], variational auto-encoder (VAE) models [35] and normaliizing flow models [36] to learn the unknown probability distribution of given image data, and entropy encode the pixel values by arithmetic coding driven by the learned probability models. These DNN models for lossless image coding have beaten the best of the traditional lossless image codecs in compression performance, reducing the lossless bit rate by about ten percent on natural color images.
@@ I Introduction
We provide a visual comparison between the traditional frameworks for noisy image compression and the proposed functionally lossless compression method in Fig. 1. In the current practice, a noisy image is either directly losslessly compressed or first denoised and then losslessly compressed. These two approaches are both sub-optimal in terms of rate-distortion metric. Specifically, direct lossless compression not only preserves the noise but also wastes bits, being detrimental to the transmission and the subsequent machine vision tasks. The cascaded approach of denoising followed by lossless compression is complex and wasteful. In contrast, the proposed fucntionally lossless compression method achieves optimal joint denoising and compression performance, and at the same time it offers higher computational efficiency and lower latency.
@@
  • We propose and implement two different deep learning based solutions respectively for two scenarios: the latent clean image is available and unavailable in the training phase.
@@
II Related Works
Image compression [8, 9, 37, 11, 12, 13, 14, 15, 16, 38, 17, 18, 20] and image denoising [39, 40, 41, 42, 43, 44, 45, 46] have been thoroughly studied by researchers in both camps of traditional image processing and modern deep learning. However, the joint image compression and denoising task has been little explored. Very few papers addressed this topic. Testolina et al. [47] investigated the integration of denoising convolutional layers in the decoder of a learning-based compression network. Ranjbar et al. [48] presented a learning-based image compression framework where image denoising and compression are performed jointly. The latent space of the image codec is organized in a scalable manner such that the clean image can be decoded from a subset of the latent space, while the noisy image is decoded from the full latent space at a higher rate. Cheng et al. [49] proposed to optimize the image compression algorithm to be noise-aware as joint denoising and compression. The key is to transform the original noisy images to noise-free bits by eliminating the undesired noise during compression, where the bits are later decompressed as clean images. Huang et al. [50] proposed an efficient end-to-end image compression network, named Noise-Adaptive ResNet VAE (NARV), aiming to handle both clean and noisy input images of different noise levels in a single noise-adaptive compression network without adding nontrivial processing time.
@@ II Related Works
III Research Problems and Methodology
III-A Problem formulation
@@
III-B Theoretical analysis of clean entropy estimation
@@
III-C Practical estimation of clean image entropy
@@
III-D Network design
The overall frameworks of the proposed supervised and weakly supervised FLLIC are illustrated in Fig. 2.
For the supervised FLLIC problem, given the noisy image $I_{n}$, we firstly encode the noisy image to the latent space for feature extraction. Then we obtain the content adaptive quantization intervals from the entropy model and apply the spatial-channel-wise quantization on the latent features. The quantized features are fed into entropy model for conditional probability estimation and for further arithmetic coding. In the decoder, the arithmetic decoded codes are fed into a decoder for noise-free image reconstruction. We minimize the distortion between the reconstructed image and the supervised clean image.
@@ III-D Network design
Figure 5: Quantitative results on the various combinations of different synthetic datasets (BSD600, Kodak24, Urban100) and different noise intensities ($\sigma$=5,10,15,20). ”FLLIC” represents the supervised FLLIC and ’FLLIC-w’ represents the weakly-supervised FLLIC.
Figure 6: Quantitative results on the real world dataset SIDD. ”FLLIC” represents the supervised FLLIC and ’FLLIC-w’ represents the weakly-supervised FLLIC.
IV Experiments
@@ IV-A Experiment setup
Datasets. Following the previous work on lossless image compression [55], we train the proposed network with Flickr2k dataset [56], which contains 2,000 high-quality images. We evaluate the trained network on three synthetic benchmark datasets generated with additive white Gaussian noise (BSD68 [57], Urban100 [58] and Kodak24 [59]). Unlike the pure image denoising task which utilizes very large noise level, we include four relatively small noise levels: $\sigma=5,10,15,20$. We also evaluate the proposed FLLIC compression method on a real-world dataset SIDD [60], which contains  30,000 noisy images from 10 scenes under different lighting conditions using five representative smartphone cameras and the corresponding generated ’noise-free’ clean images.
@@ IV-B Quantitative results
IV-C Inference time
We measure the inference time required for encoding and decoding a $512\times 512$ image on a Nvidia RTX 4090 GPU. The detailed inference time of competing methods are listed in Table I. For the cascaded method, encoding an images needs two steps: denoising and lossless compression, which need about 400 ms per step by the state-of-the-art algorithms. However, the proposed FLLIC just needs about 60 ms by one pass, which is an order of magnitude lower than the Restormer and LC-FDNet. Considering that FLLIC achieves better rate-distortion performance than the cascade scheme, it is amazing that it is also substantially ahead in inference time.
TABLE I: Inference time (ms) of cascaded method of denoising + lossless compression and the proposed FLLIC method for encoding and decoding a $512\times 512$ image on a Nvidia RTX 4090 GPU.
@@ IV-C Inference time
IV-D Ablation studies
Figure 7: Ablation results on the Kodak dataset with Gaussian noise of $\sigma$=5. ”FLLIC” represents the supervised FLLIC and ’FLLIC-w’ represents the weakly-supervised FLLIC. ”FLLIC-u” represents the FLLIC compression method with uniform quantization. ’FLLIC-w0’ represents the weakly supervised FLLIC compression network without the guidance of clean image entropy.
IV-E Limitation
@@
V Conclusion

### v2

@@ Abstract
1 Introduction
Accompanying the exciting progress of modern machine learning with deep neural networks (DNNs), many researchers have published a family of end-to-end optimized DNN image compression methods in the past five years. Most of these methods are rate-distortion optimized for lossy compression  [4, 41, 2, 5, 33, 29, 23, 7, 32, 12, 47, 19, 11, 9, 46, 24, 39, 55, 52]. By design, they cannot perform lossless or near-lossless image compression even with an unlimited bit budget. More recently, a number of research teams embark on developing DNN lossless image compression methods, aiming at minimum code length [30, 31, 42, 14, 13, 50, 51, 18]. These authors apply various deep neural networks, such as autoregressive models [44, 38], variational auto-encoder (VAE) models [21] and normaliizing flow models [22] to learn the unknown probability distribution of given image data, and entropy encode the pixel values by arithmetic coding driven by the learned probability models. These DNN models for lossless image coding have beaten the best of the traditional lossless image codecs in compression performance, reducing the lossless bit rate by about ten percent on natural color images.
@@ 1 Introduction
We provide a visual comparison between the traditional frameworks for noisy image compression and the proposed functionally lossless compression method in Fig. 1. In the current practice, a noisy image is either directly losslessly compressed or first denoised and then losslessly compressed. These two approaches are both sub-optimal in terms of rate-distortion metric. Specifically, direct lossless compression not only preserves the noise but also wastes bits, being detrimental to the transmission and the subsequent machine vision tasks. The cascaded approach of denoising followed by lossless compression is complex and wasteful. In contrast, the proposed functionally lossless compression method achieves optimal joint denoising and compression performance, and at the same time it offers higher computational efficiency and lower latency.
In practice, the clean image is never known, which means that clean images are unavailable even for training. To address this challenge, we propose to estimate the clean image entropy from the noisy observation. It is easier to estimate the entropy of the clean image than the clean image itself, as the former only involves the estimation of pmf (histogram) of K parameters vs the latter the estimation of an image signal of MN dimensions.
A key innovation of our work is the introduction of clean image entropy as a prior knowledge to guide a joint denoising and compression process. Specifically, we propose a method to estimate clean image entropy directly from noisy observations. This estimated entropy is then leveraged to adaptively predict spatial-channel-wise quantization intervals during the compression stage, with the goal of achieving a compressed latent representation whose rate closely aligns with the true entropy of the underlying noise-free image.
@@
  • We propose and implement two different deep learning based solutions respectively for two scenarios: the latent clean image is available and unavailable in the training phase, respectively.
@@
2 Related Works
Image compression [7, 32, 12, 47, 19, 11, 9, 46, 24, 39, 55] and image denoising [49, 43, 16, 26, 53, 10, 36, 48] have been thoroughly studied by researchers in both camps of traditional image processing and modern deep learning. However, the joint image compression and denoising task has been little explored. Very few papers addressed this topic. Testolina et al. [40] investigated the integration of denoising convolutional layers in the decoder of a learning-based compression network. Ranjbar et al. [35] presented a learning-based image compression framework where image denoising and compression are performed jointly. The latent space of the image codec is organized in a scalable manner such that the clean image can be decoded from a subset of the latent space, while the noisy image is decoded from the full latent space at a higher rate. Cheng et al. [6] proposed to optimize the image compression algorithm to be noise-aware as joint denoising and compression. The key is to transform the original noisy images to noise-free bits by eliminating the undesired noise during compression, where the bits are later decompressed as clean images. Huang et al. [17] proposed an efficient end-to-end image compression network, named Noise-Adaptive ResNet VAE (NARV), aiming to handle both clean and noisy input images of different noise levels in a single noise-adaptive compression network without adding nontrivial processing time.
@@ 2 Related Works
3 Research Problems and Methodology
3.1 Problem formulation
@@
3.2 Theoretical analysis of clean entropy estimation
@@
3.3 Practical estimation of clean image entropy
@@
3.4 Network design
The overall frameworks of the proposed supervised and weakly supervised FLLIC are illustrated in Fig. 2. For the supervised FLLIC problem, given the noisy image $I_{n}$, we firstly encode the noisy image to the latent space for feature extraction. Then we obtain the content adaptive quantization intervals from the entropy model and apply the spatial-channel-wise quantization on the latent features. The quantized features are fed into entropy model for conditional probability estimation and for further arithmetic coding. In the decoder, the arithmetic decoded codes are fed into a decoder for noise-free image reconstruction. We minimize the distortion between the reconstructed image and the supervised clean image.
@@ 3.4 Network design
4 Experiments
@@ 4.1 Experiment setup
Datasets. Following the previous work on lossless image compression [37], we train the proposed network with Flickr2k dataset [27], which contains 2,000 high-quality images. We evaluate the trained network on three synthetic benchmark datasets generated with additive white Gaussian noise (BSD68 [28], Urban100 [15] and Kodak24 [8]). Unlike the pure image denoising task which utilizes very large noise level, we include four relatively small noise levels: $\sigma=5,10,15$. We also evaluate the proposed FLLIC compression method on a real-world dataset SIDD [1], which contains  30,000 noisy images from 10 scenes under different lighting conditions using five representative smartphone cameras and the corresponding generated ’noise-free’ clean images.
@@ 4.2 Quantitative results
Figure 5: Quantitative results on the various combinations of different synthetic datasets (BSD600, Kodak24, Urban100) and different noise intensities ($\sigma$=5,10,15). "FLLIC" represents the supervised FLLIC and ’FLLIC-w’ represents the weakly-supervised FLLIC.
4.3 Inference time
We measure the inference time required for encoding and decoding a $512\times 512$ image on a Nvidia RTX 4090 GPU. The detailed inference time of competing methods are listed in Table 1. For the cascaded method, encoding an images needs two steps: denoising and lossless compression, which need about 400 ms per step by the state-of-the-art algorithms. However, the proposed FLLIC just needs about 60 ms by one pass, which is an order of magnitude lower than the Restormer and LC-FDNet. Considering that FLLIC achieves better rate-distortion performance than the cascade scheme, it is amazing that it is also substantially ahead in inference time.
4.4 Ablation studies
Figure 6: Quantitative results on the real world image denoising dataset SIDD medium. ’FLLIC’ represents the supervised FLLIC method and ’FLLIC-w’ represents the weakly-supervised FLLIC version.
Figure 7: Ablation results on the Kodak dataset. ’FLLIC-u’ represents the FLLIC with uniform quantization. ’FLLIC-w0’ represents the weakly supervised FLLIC without the guidance of clean image entropy.
Table 1: Inference time (ms) of cascaded method of denoising + lossless compression and the proposed FLLIC method for encoding and decoding a $512\times 512$ image on a Nvidia RTX 4090 GPU.
@@ 4.4 Ablation studies
5 Conclusion
Limitation
@@ 5 Conclusion
Broader Impact
FLLIC’s contributions to improving image quality, reducing resource usage, and enhancing efficiency and user experience make it a valuable advancement with far-reaching positive impacts across various industries and applications.
Appendix A Appendix / supplemental material
A.1 Broader Impacts
The proposed functionally lossless image compression (FLLIC) method presents several positive broader impacts:
  • Enhanced Image Quality: By jointly denoising and compressing images, FLLIC can significantly improve the perceived quality of images in various applications. This enhancement is particularly beneficial in fields such as medical imaging, where clearer images can lead to better diagnosis and treatment planning, and in remote sensing, where higher quality images improve data analysis and decision-making.
  • Reduced Storage and Bandwidth Requirements: FLLIC’s superior compression performance over mathematically lossless methods leads to lower storage and bandwidth requirements. This reduction translates into cost savings for data storage and transmission, making it particularly advantageous for industries dealing with large volumes of image data, such as digital media, cloud storage services, and scientific research.
  • Improved Efficiency in Imaging Systems: By addressing the inherent acquisition noise in digital sensors, FLLIC offers a more efficient approach to image compression. This efficiency can enhance the performance of imaging systems in various applications, from consumer electronics to professional photography, ensuring that users receive higher quality images without compromising on storage efficiency.
  • Environmental Benefits: The reduction in storage and bandwidth requirements also has positive environmental implications. Lower data storage needs mean less energy consumption in data centers, and reduced bandwidth usage can decrease the energy footprint of data transmission networks. These factors contribute to a more sustainable and eco-friendly approach to handling large-scale image data.
  • Better User Experience: In applications such as web browsing, video streaming, and online gaming, FLLIC’s ability to deliver high-quality images with reduced latency and faster loading times can significantly enhance the user experience. This improvement is especially crucial in regions with limited bandwidth or high latency networks, providing a more accessible and enjoyable digital experience.
  • Advancement in Image Processing Technology: FLLIC represents a paradigm shift in image compression, promoting further research and innovation in the field. By challenging the necessity of mathematically lossless compression and introducing a functionally lossless approach, this method opens new avenues for developing advanced image processing techniques that prioritize practical performance and quality.
Overall, FLLIC’s contributions to improving image quality, reducing resource usage, and enhancing efficiency and user experience make it a valuable advancement with far-reaching positive impacts across various industries and applications.